OpenStudy (anonymous):

Medal and Fan.. show step by step full solution: Derivative of (sin 2x) ^ (e^x)

OpenStudy (anonymous):

\[(\sin 2x)^{e ^{x}}\]

OpenStudy (anonymous):

@modphysnoob can you please explain that?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@phi

OpenStudy (phi):

try taking the ln first

OpenStudy (anonymous):

\[y=\left(\sin x\right)^{e^x}\\ \ln y=\ln (\sin x)^{e^x}\\ \ln y=e^x \ln \sin x\] Then apply some implicit differentiation.

OpenStudy (anonymous):

@SithsAndGiggles how about the 2x?

OpenStudy (anonymous):

Sorry, minor typo. Just make sure to account for it: \[\frac{d}{dx}\ln y=\frac{d}{dx}\left[e^x\ln \sin2x\right]\]

OpenStudy (anonymous):

I still need to derive sin 2x right?

OpenStudy (anonymous):

Yes. The right side will involve product rule first, then chain rule as you deal with \(\ln \sin 2x\).

OpenStudy (anonymous):

ok, i will try..